# Microphase Segregation in Polyelectrolyte Brushes

### v2

@@ Abstract
Polyelectrolyte (PE) brushes have ubiquitous applications as surface modifiers which regulate various structural and dynamic properties. Here, we apply a continuous-space self-consistent field theory to study the structural heterogeneity in PE brushes induced by competing electrostatic and hydrophobic interactions. For brushes with high grafting densities, we find a series of microphase-segregated morphologies with alternating polymer-rich and polymer-poor layers in the direction normal to the substrate. The transitions between multi-layer morphologies with consecutive numbers of condensed layers as well as the melting transition to the fully swollen brush are all discontinuous. We also elucidate that the segregated layers are formed by different subpopulations of all chains, significantly different from the scenario of the pearl-necklace structure formed by a single PE in poor solvents. Furthermore, we bridge the microstructure of multi-layer brushes to experimentally measurable reflectivity spectra, where the oscillation period and amplitude in the spectra are shown to be very sensitive to the number of layers and the sharpness of the polymer-solvent interface. The multi-layer morphology predicted by our theory is in good agreement with the lamellae structure experimentally observed at hydrated Nafion film interfaces.
[…]
@@ Introduction
Uncovering the microstructure of PE brushes using state-of-the-art experimental techniques remains a great challenge. Computational methods provide an alternative solution. Great efforts have been made to theoretically model PE brushes since the pioneering work of de Gennes and Pincus 29, 30. Built upon the Alexander—de Gennes model, Pincus developed a scaling theory which utilized a force balance between the repulsive osmotic pressure due to the translational entropy of ions and the elastic attraction due to chain deformation. Zhulina and coworkers presented a diagram of states including the osmotic, salted and neutral brush regimes 31, 32, 33. In addition, lattice self-consistent field theory (commonly known as the Scheutsjen—Fleer model) has been widely applied to describe synthetic and protein-inspired PE brushes 34, 35, 36, 37, 38. However, the Scheutsjen—Fleer model constrains each monomer to a lattice site, which limits its flexibility in describing complex heterogeneous morphologies. Furthermore, classical density functional theory (CDFT) has also been used to investigate PE brush morphologies. The predicted scaling relationship between the brush height and salt concentration is in agreement with scaling theory and simulations 39, 40. However, CDFT largely relies on the preassumed functional form of the electrostatic correlation and intra-chain correlation, where improper choices can lead to non-physical results 41, 42, 43.
@@
The coexistence of local segregation and long-range stretching is an important conformational feature which widely exists in various PE systems 44, 45, 46, 47. A well-known example is the pearl-necklace structure formed by a single PE in poor solvents 48, 49, 50, 51. It is thus natural to ask how this structure changes when PEs are collectively tethered to a surface. The study of structural heterogeneity with local segregation is critical for understanding the morphology, surface properties and interactions of PE brushes in order to fully realize their potential as surface regulators. In an earlier work, Duan et al. predicted that structural heterogeneity is preferred in the lateral direction for brushes with lower grafting densities, whereas brushes with higher grafting densities favor heterogeneous morphologies in the normal direction 52. In this work, we apply continuous-space self-consistent field theory (SCFT) to systematically investigate the effects of the competing electrostatic and hydrophobic interactions on the structural heterogeneity of PE brushes with high grafting densities. We find cascade transitions between a series of microphase-segregated morphologies with alternating polymer-rich and polymer-poor layers in the direction normal to the substrates. The molecular picture underlying these multi-layer morphologies is also discussed. Furthermore, we examine the evolution of scattering spectra accompanying the morphological changes.
@@
As shown in Fig. 1, we consider brushes composed of polyelectrolyte (PE) chains with one end grafted on a substrate and immersed in a solution of $n_{s}$ solvent molecules and $n_{\pm}$ mobile ions. The system is treated as a semicanonical ensemble with a fixed number of PE chains $n_{p}$ whereas solvent and mobile ions are connected to a bulk salt solution of ion concentration $c_{\pm}^{b}$ that maintains the chemical potentials of the solvent $\mu_{s}$ and ions $\mu_{\pm}$. PEs are assumed to be Gaussian chains of $N$ total segments with Kuhn length $b$. We adopt the smear charge model with the backbone charge density $\alpha$. Mobile ions are considered to be point charges with valency $z_{\pm}$.
@@
$\Xi=~{}$ $\frac{1}{n_{p}!\nu_{p}^{Nn_{p}}}\prod_{\gamma}\sum_{n_{\gamma}=0}^{\infty}\frac{e^{\mu_{\gamma}n_{\gamma}}}{n_{\gamma}!v_{\gamma}^{n_{\gamma}}}\prod_{i=1}^{n_{p}}\int\mathcal{D}\{\mathbf{R}_{i}\}$ (1)
$\prod_{j=1}^{n_{\gamma}}\int\mathrm{d}\mathbf{r}_{\gamma,j}\ \exp\bigg{\{}-\mathcal{H}\bigg{\}}\prod_{\mathbf{r}}\delta[\hat{\phi}_{p}(\mathbf{r})+\hat{\phi}_{s}(\mathbf{r})-1]~{},$
where small molecules are denoted $\gamma=s,\pm$ for solvent and ions, respectively. $\nu_{p}$ is the volume of PE segments, whereas $\nu_{\gamma}$ is the volume of small molecules. For simplicity, we assume $\nu_{p}=\nu_{s}=\nu$. $\mathcal{D}\{\mathbf{R}_{i}\}$ denotes the integration over all chain configurations for each chain $i$. $\hat{\phi}_{p}(\mathbf{r})$ and $\hat{\phi}_{s}(\mathbf{r})$ are the local instantaneous volume fractions of the PE and solvent, respectively. The $\delta$ function at the end of Eq. 1 accounts for the incompressibility.
@@
which contains the elastic energy governed by Gaussian chain statistics, the short-range hydrophobic interactions between polymer segments and solvent molecules, and the long-range Coulomb interactions between all charged species. The hydrophobic interaction is manifested by the Flory—Huggins $\chi$ parameter. $\hat{\rho}_{c}(\mathbf{r})=z_{+}\hat{c}_{+}(\mathbf{r})-z_{-}\hat{c}_{-}(\mathbf{r})+\alpha\hat{\phi}_{p}(\mathbf{r})/\nu$ is the local charge density with $\hat{c}_{\pm}(\mathbf{r})$ as the instantaneous number density of ions. $C(\mathbf{r},\mathbf{r}^{\prime})$ is the Coulomb operator satisfying $-\nabla\cdot\big{[}\epsilon(\mathbf{r})\nabla C(\mathbf{r},\mathbf{r}^{\prime})\big{]}=\delta(\mathbf{r}-\mathbf{r}{{}^{\prime}})$. $\epsilon(\mathbf{r})=kT\epsilon_{0}\epsilon_{r}(\mathbf{r})/e^{2}$ is the scaled permittivity, where $\epsilon_{0}$ is the vacuum permittivity, $e$ is the elementary charge, and $\epsilon_{r}(\mathbf{r})$ is the local dielectric constant dependent on the local composition of the system.
We follow the standard self-consistent field procedure 53 by first decoupling the interacting system into non-interacting PE chains and ions in fluctuating fields using the Hubbard—Stratonovich transformation and identity transforms. We then employ the saddle-point approximation to simplify the evaluation of the functional integral over the fluctuating fields. These give rise to the following self-consistent equations for polymer density $\phi_{p}(\mathbf{r})$, conjugate fields $w_{p}(\mathbf{r})$ and $w_{s}(\mathbf{r})$, electrostatic potential $\psi(\mathbf{r})$ and ion concentrations $c_{\pm}(\mathbf{r})$:
@@
$-\frac{\partial\epsilon(\mathbf{r})}{\partial\phi_{p}}\nu\lvert\nabla\psi(\mathbf{r})\rvert^{2}-\chi\phi_{p}(\mathbf{r})$
$\phi_{p}(\mathbf{r})$ $=\frac{n_{p}}{Q_{p}}\int_{0}^{N}\mathrm{ds}\ q(\mathbf{r};\mathrm{s})q_{c}(\mathbf{r};\mathrm{s})$ (3b)
@@
$-\nabla\cdot\big{(}\epsilon(\mathbf{r})\nabla\psi(\mathbf{r})\big{)}$ $=z_{+}c_{+}(\mathbf{r})-z_{-}c_{-}(\mathbf{r})+\frac{\alpha}{\nu}\phi_{p}(\mathbf{r})~{},$ (3d)
where $c_{\pm}=\lambda_{\pm}\exp(\mp z_{\pm}\psi)$ is the ion concentration with $\lambda_{\pm}=e^{\mu_{\pm}}/\nu_{\pm}$ the fugacity of the ions determined by the bulk ion concentration $c^{b}_{\pm}$. $Q_{p}=\nu^{-1}\int\mathrm{d}\mathbf{r}\ q(\mathbf{r};s)$ is the single-chain partition function, where $q(\mathbf{r};s)$ is the chain propagator satisfying the modified diffusion equation:
$\frac{\partial}{\partial s}q(\mathbf{r};\mathrm{s})=\frac{b^{2}}{6}$ $\nabla^{2}q(\mathbf{r};s)-w_{p}(\mathbf{r})q(\mathbf{r};s)~{}.$ (4)
@@
$F=$ $-\sigma\ln Q_{p}-e^{\mu_{s}}Q_{s}$
$+\frac{1}{\nu}\int\mathrm{dz}\ \bigg{(}\chi\phi_{p}(1-\phi_{p})-w_{p}\phi_{p}-w_{s}(1-\phi_{p})\bigg{)}$
$+\int\mathrm{dz}\ \bigg{[}-\frac{\epsilon}{2}\bigg{(}\frac{\mathrm{d}\psi}{\mathrm{dz}}\bigg{)}^{2}+\frac{\alpha}{\nu}\phi_{p}\psi-c_{+}-c_{-}+c_{+}^{b}+c_{-}^{b}\bigg{]}~{}.$ (5)
@@
$\sigma=n_{p}/A$ is the grafting density of chains on the substrate, where $A$ is the total area of the substrate. $Q_{s}=\nu^{-1}\int\mathrm{dz}\ \exp(-w_{s})$ is the solvent partition function. The density of PEs $\phi_{p}$ and the gradient of the electrostatic potential $\psi$ are set to zero on the boundaries of the system.
@@
Eq. 3 is solved iteratively until self-consistency is achieved and the equilibrium PE brush density distribution, electrostatic potential, and ion distribution can be obtained. The modified diffusion equation (Eq. 4) is solved by using the Crank—Nicolson algorithm whereas the Poisson—Boltzmann equation (Eq. 3d) is solved using an iterative, centered finite difference scheme 54. In contrast to the Scheutsjen—Fleer (SF) model with the lattice constraint 35, 36, the differential equations in our work are solved in the continuous space, which decouple the numerical discretization from the physical lattice. This improves both the accuracy and flexibility of the calculations, enabling us to explore morphologies with more complex heterogeneity 55. Furthermore, our theory can not only be easily generalized to polyelectrolytes with various chain architectures, copolymer composition and charge patterns, but also be straightfowardly implemented to biomacromolecules like disordered protein brushes 56, 57.
Results and Discussion
In this work, we focus on the competing interactions between electrostatic repulsion and hydrophobic attraction on the morphology of PE brushes. These two interactions are controlled by the backbone charge density $\alpha$ and the Flory—Huggins $\chi$ parameter, respectively. We consider dense brushes with grafting density $\sigma=0.02\ \mathrm{nm}^{-2}$ immersed in a monovalent salt ($z_{\pm}=1$) solution with bulk concentration $c_{\pm}^{b}=10$ mM. For simplicity, the dielectric constant of the system is assumed to be uniform and set to be $\epsilon_{r}=80$, the value of water. The temperature is set at 293 K, yielding a Bjerrum length $l_{B}=e^{2}/4\pi kT\epsilon_{0}\epsilon_{r}$ of 0.7 nm, and a Debye screening length $\kappa_{D}^{-1}=[4\pi l_{B}(z_{+}c_{+}^{b}+z_{-}c_{-}^{b})]^{-1/2}$ of 3 nm. Each constituent polymer has a chain length of 200 segments with the Kuhn length $b=1.0$ nm.
Evolution of Multi-layer Morphologies
It is well-known that a single PE chain in poor solvents can form a series of pearl-necklace structures induced by the competition between long-range electrostatic repulsion and short-range hydrophobic attraction 49, 50, 58. Similar heterogeneous structures with local polymer segregation also exist in PE brushes. Fig. 2 shows that, as the hydrophobicity $\chi$ decreases with the charge density fixed at $\alpha=0.4$, a series of microphase-segregated multi-layer morphologies appear. When hydrophobic attraction is dominant ($\chi=2.0$), all the polymers are compressed into a single condensed layer to minimize the contact with the solvent. As $\chi$ decreases, hydrophobic attraction is reduced—equivalently, the relative importance of electrostatic interaction is enhanced. PE brushes form increasing numbers of condensed layers (e.g., from two at $\chi=1.70$ to four at $\chi=1.30$) to increase the average separation distance between charges. The widths of these additional layers reflect the size of an electrostatic blob, approximately the Debye screening length $\kappa_{D}^{-1}$. For a given $\alpha$ and $\chi$, the maximum polymer density of each condensed layer is almost the same, which is determined by the maximum local charge density that can be maintained. The distance between two adjacent condensed layers is determined by balancing the electrostatic repulsion between the layers and the elastic energy due to chain stretching. Our theoretical prediction of these multi-layer morphologies is consistent with the experimental findings by Dura et al. and Randall et al. of the lamellae structure composed of alternating layers of ionomer Nafion and water at the interface between the substrate and the hydrated Nafion film 27, 28.
As $\chi$ further decreases to $\chi=1.20$ (Fig. 2), an additional swollen layer of lower polymer density is melted from the inner condensed layers and extends far into the solution. The coexistence between inner condensed layers and an outer swollen layer predicted by our theory is in good agreement with the experimental observations via cryo-TEM images of PSS-grafted particles by Samokhina et al. This also explains the non-trivial neutron reflectivity spectra reported by Reinhardt et al. in the study of neutral and charged brushes 25, 59. Finally, when the hydrophobicity becomes low, e.g. $\chi=1.00$ (Fig. 2), PE brushes are fully swollen, recovering the scenario of the Alexander—de Gennes brush in the salted regime.
@@
By systematically including the electrostatics into the polymeric self-consistent field theory, we are able to capture the interplay between the charge repulsion and hydrophobic attraction on the structural heterogeneity in PE brushes. Fig. 4 shows the evolution of multi-layer morphologies for different backbone charge densities $\alpha$. The morphology is characterized by the brush height $H$, a macroscopic measurable property which can be obtained experimentally using methods like ellipsometry and AFM. To facilitate the characterization of the multi-layer structures, $H$ in the current work is defined as the distance at which the polymer density reduces to a threshold $\phi_{p}^{*}=1\times 10^{-6}$. Another commonly adopted definition of brush height based on the Gibbs dividing surface 60, 61 is not sensitive to structural heterogeneity.
As shown in Fig. 4, PE brushes exhibit richer morphological behaviors with more accessible multi-layer structures as $\alpha$ increases. Due to the absence of the long-range electrostatic repulsion, charge neutral brushes ($\alpha=0$) only exhibit a continuous transition from a fully condensed morphology to a fully swollen morphology. As $\alpha$ increases to 0.2, it is interesting to note that although an outer swollen layer emerges in the pre-swollen process, the transition remains continuous. This is consistent with early theoretical work invoking the assumption of local charge neutrality 62. For larger backbone charge densities (e.g., $\alpha\geq 0.3$), morphologies with multiple condensed layers appear. As $\alpha$ increases up to $0.4$, morphologies with two, three and four condensed layers are all accessible. The stronger electrostatic repulsion provides a larger driving force to overcome the energetic penalty associated with the increase in surface area when a new condensed layer is generated.

### v3

@@ Abstract
Polyelectrolyte (PE) brushes have ubiquitous applications as surface modifiers which regulate various structural and dynamic properties. Here, we apply a continuous-space self-consistent field theory to study the structural heterogeneity in PE brushes induced by competing electrostatic and hydrophobic interactions. For brushes with high grafting densities, we find a series of microphase-segregated morphologies with alternating polymer-rich and polymer-poor layers in the direction normal to the substrate. The transitions between multi-layer morphologies with consecutive numbers of condensed layers as well as the melting transition to the fully swollen brush are all discontinuous. We also elucidate that the segregated layers are formed by different subpopulations of all chains, significantly different from the scenario of the pearl-necklace structure formed by a single PE in poor solvents. Furthermore, we bridge the microstructure of multi-layer brushes to experimentally measurable reflectivity spectra, where the oscillation period and amplitude in the spectra are shown to be very sensitive to the number of layers and the sharpness of the polymer-solvent interface. The multi-layer morphology predicted by our theory is in good agreement with the lamellae structure experimentally observed at hydrated Nafion film interfaces. {tocentry} [Uncaptioned image]
@@ Introduction
Uncovering the microstructure of PE brushes using state-of-the-art experimental techniques remains a great challenge. Computational methods provide an alternative solution 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48. Great efforts have been made to theoretically model PE brushes since the pioneering work of de Gennes and Pincus 29, 30. Built upon the Alexander—de Gennes model, Pincus developed a scaling theory which utilized a force balance between the repulsive osmotic pressure due to the translational entropy of ions and the elastic attraction due to chain deformation. Zhulina and coworkers presented a diagram of states including the osmotic, salted and neutral brush regimes 31, 32, 33. In addition, lattice self-consistent field theory (commonly known as the Scheutjens—Fleer model) has been widely applied to describe synthetic and protein-inspired PE brushes 34, 35, 36, 37, 38. However, the Scheutjens—Fleer model constrains each monomer to a lattice site, which limits its flexibility in describing complex heterogeneous morphologies. Furthermore, classical density functional theory (CDFT) has also been used to investigate PE brush morphologies. The predicted scaling relationship between the brush height and salt concentration is in agreement with scaling theory and simulations 39, 40. However, CDFT largely relies on the preassumed functional form of the electrostatic correlation and intra-chain correlation, where improper choices can lead to non-physical results 41, 42, 43.
@@
The coexistence of local segregation and long-range stretching is an important conformational feature of PE49, 50, 51, 52. This may lead to the emergence of microphase separation in a variety of PE systems 53, 54, 55. A well-known example is the pearl-necklace structure formed by a single PE in poor solvents 56, 57, 58, 59. Recently, Liu et al. developed a variational approach to control the center-of-mass of polymers 60. Incorporating this method into the self-consistent field theory (SCFT), they are able to predict the pearl-necklace structure formed by a single PE. The predicted polymer conformation, scaling behavior and cascade transition is in good agreement with the classical work by Dobrynin, Rubinstein, and Obukhov 56. It is thus natural to ask how this structure changes when PEs are collectively tethered to a surface. The study of structural heterogeneity with local segregation is critical for understanding the morphology, surface properties and interactions of PE brushes in order to fully realize their potential as surface regulators. In an earlier work, Duan et al. predicted that structural heterogeneity is preferred in the lateral direction for brushes with lower grafting densities, whereas brushes with higher grafting densities favor heterogeneous morphologies in the normal direction 61. In this work, we apply continuous-space SCFT to systematically investigate the effects of the competing electrostatic and hydrophobic interactions on the structural heterogeneity of PE brushes with high grafting densities. We find cascade transitions between a series of microphase-segregated morphologies with alternating polymer-rich and polymer-poor layers in the direction normal to the substrates. The molecular picture underlying these multi-layer morphologies is also discussed. Furthermore, we examine the evolution of scattering spectra accompanying the morphological changes.
@@
As shown in Fig. 1, we consider brushes composed of polyelectrolyte (PE) chains with one end grafted on a substrate and immersed in a solution of $n_{s}$ solvent molecules and $n_{\pm}$ mobile ions. The system is treated as a semicanonical ensemble with a fixed number of PE chains $n_{p}$ whereas solvent and mobile ions are connected to a bulk salt solution of ion concentration $c_{\pm}^{b}$ that maintains the chemical potentials of the solvent $\mu_{s}$ and ions $\mu_{\pm}$. PEs are assumed to be Gaussian chains of $N$ total segments with Kuhn length $b$. We adopt the smear charge model with the backbone charge density $\alpha$. Mobile ions are considered to have valency $z_{\pm}$.
@@
$\Xi=~{}$ $\frac{1}{n_{p}!\nu_{p}^{Nn_{p}}}\prod_{\gamma}\sum_{n_{\gamma}=0}^{\infty}\frac{e^{\mu_{\gamma}n_{\gamma}}}{n_{\gamma}!v_{\gamma}^{n_{\gamma}}}$ (1)
$\prod_{i=1}^{n_{p}}\int\mathcal{D}\{\mathbf{R}_{i}\}\prod_{j=1}^{n_{\gamma}}\int\mathrm{d}\mathbf{r}_{\gamma,j}\ \exp\bigg{\{}-\mathcal{H}\bigg{\}}$
$\prod_{\mathbf{r}}\delta[\hat{\phi}_{p}(\mathbf{r})+\hat{\phi}_{s}(\mathbf{r})-1]~{},$
where small molecules are denoted $\gamma=s,\pm$ for solvent and ions, respectively. $\nu_{p}$ is the volume of PE segments, whereas $\nu_{\gamma}$ is the volume of small molecules. For simplicity, we assume $\nu_{p}=\nu_{s}=\nu$. $\mathcal{D}\{\mathbf{R}_{i}\}$ denotes the integration over all chain configurations for each chain $i$. $\hat{\phi}_{p}(\mathbf{r})$ and $\hat{\phi}_{s}(\mathbf{r})$ are the local instantaneous volume fractions of the PE and solvent, respectively. The $\delta$ function at the end of Eq. 1 accounts for the incompressibility. We only consider the contributions from polymers and solvent in the incompressibility constraint. For the parameter space studied in this work, the volume fractions of ions in the brushes are very low and can thus be neglected.
@@
which contains the elastic energy governed by Gaussian chain statistics, the short-range hydrophobic interactions between polymer segments and solvent molecules, and the long-range Coulomb interactions between all charged species. The hydrophobic interaction is manifested by the Flory—Huggins $\chi$ parameter. $\hat{\rho}_{c}(\mathbf{r})=z_{+}\hat{c}_{+}(\mathbf{r})-z_{-}\hat{c}_{-}(\mathbf{r})+\alpha\hat{\phi}_{p}(\mathbf{r})/\nu$ is the local charge density with $\hat{c}_{\pm}(\mathbf{r})$ as the instantaneous number density of ions. $C(\mathbf{r},\mathbf{r}^{\prime})$ is the Coulomb operator satisfying $-\nabla\cdot\big{[}\epsilon(\mathbf{r})\nabla C(\mathbf{r},\mathbf{r}^{\prime})\big{]}=\delta(\mathbf{r}-\mathbf{r}{{}^{\prime}})$. $\epsilon(\mathbf{r})=kT\epsilon_{0}\epsilon_{r}(\mathbf{r})/e^{2}$ is the scaled permittivity, where $\epsilon_{0}$ is the vacuum permittivity, $e$ is the elementary charge, and $\epsilon_{r}(\mathbf{r})$ is the local dielectric constant dependent on the local composition of the system 62, 63, 64, 65.
We follow the standard self-consistent field procedure 66 by first decoupling the interacting system into non-interacting PE chains and ions in fluctuating fields using the Hubbard—Stratonovich transformation and identity transforms. We then employ the saddle-point approximation to simplify the evaluation of the functional integral over the fluctuating fields. These give rise to the following self-consistent equations for polymer density $\phi_{p}(\mathbf{r})$, conjugate fields $w_{p}(\mathbf{r})$ and $w_{s}(\mathbf{r})$, electrostatic potential $\psi(\mathbf{r})$ and ion concentrations $c_{\pm}(\mathbf{r})$:
@@
$\quad\quad\quad\quad\quad\quad\quad-\frac{\partial\epsilon(\mathbf{r})}{\partial\phi_{p}}\nu\lvert\nabla\psi(\mathbf{r})\rvert^{2}-\chi\phi_{p}(\mathbf{r})$
$\phi_{p}(\mathbf{r})=~{}\frac{n_{p}}{Q_{p}}\int_{0}^{N}\mathrm{ds}\ q(\mathbf{r};\mathrm{s})q_{c}(\mathbf{r};\mathrm{s})$ (3b)
@@
$-\nabla\cdot\big{(}\epsilon(\mathbf{r})\nabla\psi(\mathbf{r})\big{)}=$ (3d)
$\quad\quad\quad\quad\quad\quad\quad z_{+}c_{+}(\mathbf{r})-z_{-}c_{-}(\mathbf{r})+\frac{\alpha}{\nu}\phi_{p}(\mathbf{r})~{},$
where $c_{\pm}=\lambda_{\pm}\exp(\mp z_{\pm}\psi)$ is the ion concentration with $\lambda_{\pm}=e^{\mu_{\pm}}/\nu_{\pm}$ the fugacity of the ions determined by the bulk ion concentration $c^{b}_{\pm}$. $Q_{p}=\nu^{-1}\int\mathrm{d}\mathbf{r}\ q(\mathbf{r};\mathrm{s})$ is the single-chain partition function, where $q(\mathbf{r};\mathrm{s})$ is the chain propagator satisfying the modified diffusion equation:
$\frac{\partial}{\partial\mathrm{s}}q(\mathbf{r};\mathrm{s})=\frac{b^{2}}{6}$ $\nabla^{2}q(\mathbf{r};\mathrm{s})-w_{p}(\mathbf{r})q(\mathbf{r};\mathrm{s})~{}.$ (4)
@@
$F=$ $-\sigma\ln Q_{p}-e^{\mu_{s}}Q_{s}+\frac{1}{\nu}\int\mathrm{dz}\ \Big{(}\chi\phi_{p}(1-\phi_{p})$
-- $w_{p}\phi_{p}-w_{s}(1-\phi_{p})\Big{)}+\int\mathrm{dz}\ \Big{(}\frac{\alpha}{\nu}\phi_{p}\psi$ (5)
-- $\frac{\epsilon}{2}\bigg{(}\frac{\mathrm{d}\psi}{\mathrm{dz}}\bigg{)}^{2}-c_{+}-c_{-}+c_{+}^{b}+c_{-}^{b}\Big{)}~{}.$
@@
$\sigma=n_{p}/A$ is the grafting density of chains on the substrate, where $A$ is the total area of the substrate. $Q_{s}=\nu^{-1}\int\mathrm{dz}\ \exp(-w_{s})$ is the solvent partition function. Here we consider the substrate to be impenetrable and charge neutral, which is reflected by setting the boundary conditions $\phi_{p}=0$ and $\mathrm{d}\psi/\mathrm{dz}=0$ at the surface ($z=0$).
@@
Eq. 3 is solved iteratively until self-consistency is achieved and the equilibrium PE brush density distribution, electrostatic potential, and ion distribution can be obtained. The modified diffusion equation (Eq. 4) is solved by using the Crank—Nicolson algorithm whereas the Poisson—Boltzmann equation (Eq. 3d) is solved using an iterative, centered finite difference scheme 67. In contrast to the Scheutjens—Fleer (SF) model with the lattice constraint 35, 36, the differential equations in our work are solved in the continuous space, which decouples the numerical discretization from the physical lattice. This improves both the accuracy and flexibility of the calculations, enabling us to explore morphologies with more complex heterogeneity 68. Furthermore, our theory can not only be easily generalized to polyelectrolytes with various chain architectures, copolymer composition and charge patterns, but also be straightfowardly implemented to biomacromolecules like disordered protein brushes 69, 70. In the current version of our SCFT, the electrostatics is treated at the mean-field level. The ion–ion correlations can be included by involving the Gaussian fluctuation theory into the treatment of the electrostatic fields 71, 61. We also note that, compared to methods like molecular dynamics simulations, our current theory is limited to modeling equilibrium morphologies and cannot capture the kinetics of their transitions and other dynamic properties.
@@
Results and Discussion
In this work, we focus on the competing interactions between electrostatic repulsion and hydrophobic attraction on the morphology of PE brushes. These two interactions are controlled by the backbone charge density $\alpha$ and the Flory—Huggins $\chi$ parameter, respectively. We consider PE brushes with grafting density $\sigma=0.02\ \mathrm{nm}^{-2}$ immersed in a monovalent salt ($z_{\pm}=1$) solution with bulk concentration $c_{\pm}^{b}=10$ mM. The temperature is set at 293 K, yielding a Bjerrum length $l_{B}=e^{2}/4\pi kT\epsilon_{0}\epsilon_{r}$ of 0.7 nm, and a Debye screening length $\kappa_{D}^{-1}=[4\pi l_{B}(z_{+}c_{+}^{b}+z_{-}c_{-}^{b})]^{-1/2}$ of 3 nm. Each constituent polymer has a chain length of 200 segments with the Kuhn length $b=1.0$ nm. For this set of parameters, the brushes studied in this work are within the dense brush regime. For example, with $\alpha=0.4$, $R_{g}^{2}\sigma\sim(\alpha Nb)^{2}\sigma=128\gg 1$, where $R_{g}$ is the radius of gyration of a polyelectrolyte.
Evolution of Multi-layer
Morphologies
It is well-known that a single PE chain in poor solvents can form a series of pearl-necklace structures induced by the competition between long-range electrostatic repulsion and short-range hydrophobic attraction 57, 72. Similar heterogeneous structures with local polymer segregation also exist in PE brushes. Fig. 2 shows that, as the hydrophobicity $\chi$ decreases with the charge density fixed at $\alpha=0.4$, a series of microphase-segregated multi-layer morphologies appear. When hydrophobic attraction is dominant ($\chi=2.0$), all the polymers are compressed into a single condensed layer to minimize the contact with the solvent. As $\chi$ decreases, hydrophobic attraction is reduced—equivalently, the relative importance of electrostatic interaction is enhanced. PE brushes form increasing numbers of condensed layers (e.g., from two at $\chi=1.70$ to four at $\chi=1.30$) to increase the average separation distance between charges. The widths of these additional layers reflect the size of an electrostatic blob, approximately the Debye screening length $\kappa_{D}^{-1}$. For a given $\alpha$ and $\chi$, the maximum polymer density of each condensed layer is almost the same, which is determined by the maximum local charge density that can be maintained. The distance between two adjacent condensed layers is determined by balancing the electrostatic repulsion between the layers and the elastic energy due to chain stretching. Our theoretical prediction of these multi-layer morphologies is consistent with the experimental findings by Dura et al. and Randall et al. of the lamellae structure composed of alternating layers of ionomer Nafion and water at the interface between the substrate and the hydrated Nafion film 27, 28.
As $\chi$ further decreases to $\chi=1.20$ (Fig. 2), an additional swollen layer of lower polymer density is melted from the inner condensed layers and extends far into the solution. The coexistence between inner condensed layers and an outer swollen layer predicted by our theory is in good agreement with the experimental observations via cryo-TEM images of PSS-grafted particles by Samokhina et al. 24 This also explains the non-trivial neutron reflectivity spectra reported by Reinhardt et al. in the study of neutral and charged brushes, where coexistent condensed and swollen layers were observed by fitting the spectra data.25, 73. Similar coexisting morphologies were also observed in neurofilament-derived protein brushes at intermediate salt concentrations. Using the same continuous-space SCFT as used in this work, we have previously shown quantitative agreement with experimentally measured brush heights 69, 70. Finally, when the hydrophobicity becomes low, e.g. $\chi=1.00$ (Fig. 2), PE brushes are fully swollen, recovering the scenario of the Alexander—de Gennes brush in the salted regime 32, 74. Consistent with previous work 31, 33, the brush height $H$ decreases as salt concentration $c_{\pm}^{b}$ increases, approximately following the scaling of $H\sim(c_{\pm}^{b})^{-1/3}$ (see Fig. S1 in the SI).
Figure 3: Electrostatic potential profiles $\psi$ corresponding to the multi-layer morphologies shown in Fig. 2.
[…]
Fig. 3 plots representative electrostatic potential profiles $\psi$ corresponding to the multi-layer morphologies shown in Fig. 2. The distributions of cations and anions are also shown in Fig. S2 in the SI. The multi-peak features are clearly observed in the electrostatic potential profiles, similar to their corresponding polymer density distributions. $\psi$ exhibits peaks corresponding to each polymer-rich layer and decays in each solvent-rich layer. The peak value of $\psi$ for each polymer-rich layer decreases as the layers become further away from the grafting surface. Furthermore, as shown in Fig. 3, the appearance of a “shoulder” at the outer edge of the $\psi$ profile signifies the existence of the additional swollen layer. Therefore, the multi-layer morphology of the PE brushes can be directly detected, if the average electrostatic potential in each layer can be measured experimentally.
By systematically including the electrostatics into the polymeric self-consistent field theory, we are able to capture the interplay between the charge repulsion and hydrophobic attraction on the structural heterogeneity in PE brushes. Fig. 5 shows the evolution of multi-layer morphologies for different backbone charge densities $\alpha$. The morphology is characterized by the brush height $H$, a macroscopic measurable property which can be obtained experimentally using methods like ellipsometry and AFM. To facilitate the characterization of the multi-layer structures, $H$ in the current work is defined as the distance at which the polymer density reduces to a threshold $\phi_{p}^{*}=1\times 10^{-6}$. Another commonly adopted definition of brush height based on the Gibbs dividing surface 75, 76 is not sensitive to structural heterogeneity.
@@
As shown in Fig. 5, PE brushes exhibit richer morphological behaviors with more accessible multi-layer structures as $\alpha$ increases. Due to the absence of the long-range electrostatic repulsion, charge neutral brushes ($\alpha=0$) only exhibit a continuous transition from a fully condensed morphology to a fully swollen morphology. As $\alpha$ increases to 0.2, it is interesting to note that although an outer swollen layer emerges in the pre-swollen process, the transition remains continuous. This is consistent with early theoretical work invoking the assumption of local charge neutrality 77. For larger backbone charge densities (e.g., $\alpha\geq 0.3$), morphologies with multiple condensed layers appear. As $\alpha$ increases up to $0.4$, morphologies with two, three and four condensed layers are all accessible. If the brushes are in the salted regime 32, a similar trend can also be expected as decreasing the bulk salt concentration $c_{\pm}^{b}$ (see Fig. S1 in the SI). The stronger electrostatic repulsion provides a larger driving force to overcome the energetic penalty associated with the increase in surface area when a new condensed layer is generated.
@@
As an initial attempt to study the multi-layer morphologies in PE brushes, we assume the same dielectric constants for polymers and solvent. However, as indicated by Eq. 3, our theory is general for non-uniform dielectric environments. In many systems, polymers may have a much lower dielectric constant than solvent, which leads to a stronger screening effect on the electrostatic repulsion within the layers. This will shift the transitions between morphologies with consecutive numbers of layers to lower values of $\chi$, thus suppressing the formation of multi-layer morphologies (see Fig. S3 in the SI). Therefore, lowering the dielectric constant of polymers has the similar effect as decreasing $\alpha$. Furthermore, the current study is focused on dense brushes where $R_{g}^{2}\sigma\gg 1$. We assume that the formation of multi-layers via microphase segregation in the normal direction to the surface is favorable in the dense brush regime. We note that as the grafting density decreases, lateral microphase segregation in the form of cylindrical aggregates, pinned micelles, etc. may also occur. We expect that laterally segregated morphologies will be energetically favorable in the regime of lower grafting density 23.